OpenStudy (anonymous):

∫ x^3 e^-x dx integrate using the appropriate technique of integration.

OpenStudy (anonymous):

Use some integration by parts on that guy; let u = x^3 and v'=e^-x.

OpenStudy (anonymous):

ok.. u=x^3 v=e^-x u'=3x^2 v'= ...????

OpenStudy (anonymous):

No, dv/dx = e^-x. You can't set v without first setting v', when doing integration by parts.

OpenStudy (anonymous):

You're going to end up doing that type of integration 2 or 3 times if I'm not mistaken; unless you're familiar with tabular integration.

OpenStudy (anonymous):

∫ U dV = UV -∫ V dU ???? thats right..???

OpenStudy (anonymous):

Yes, once you get that second ∫v du you're going to get another ugly one, so do integration parts on that one too.

OpenStudy (anonymous):

ohhhh... confius..